# Chen, Simchi-Levi, and Wang

### v2

@@ 3 Preliminaries: Inference under A Fixed Policy
Despite of the difference in nature, $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ are actually providing the same point estimation of the value function ${\bm{V}}$. More formally, we have the following proposition.
@@
Let $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ be the respective output of Algorithm 1 and Algorithm 2 given the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ sampled under policy $\pi$. Assume all states are visited at least once in $\tau$, then it holds $\widehat{\bm{V}}^{\textnormal{MB}}=\widehat{\bm{V}}^{\textnormal{TD}}$.
Proof 3.2
Proof of Proposition 3.1. Let $N(s^{i})>0$ be the number of visitations to state $s^{i}\in{\mathcal{S}}$ in $\tau$. Then the optimization problem in Algorithm 2 can be rewritten as
$\min_{\widehat{\bm{V}}}\left\|{\operatorname{Diag}(N(s^{1}),N(s^{2}),\ldots,N(s^{K}))\cdot\left({\widehat{\bm{V}}-\left({\widehat{\bm{r}}+\gamma\widehat{\bm{P}}\widehat{\bm{V}}}\right)}\right)}\right\|_{2}^{2}\,,$ (4)
where $\widehat{\bm{r}}\in\mathbb{R}^{K}$ is the empirical average reward with $\widehat{\bm{r}}(s)=\sum_{i=1}^{T}\mathbf{1}[s_{i}=s]r(s_{i},a_{i})/N(s)$ for $s\in{\mathcal{S}}$ and $\widehat{\bm{P}}\in\mathbb{R}^{K\times K}$ is the empirical transition matrix with $\widehat{\bm{P}}(s,s^{\prime})=\sum_{i=1}^{T}\mathbf{1}[s_{i}=s,s_{i+1}=s^{\prime}]/N(s)$ for all $s,s^{\prime}\in{\mathcal{S}}$.
[…]
@@
We have now demonstrated the equivalence between the two estimators within finite-space and finite action MDPs without function approximations, indicating that they share identical theoretical properties. Consequently, establishing a property for one estimator inherently confirms it for the other. It is important to emphasize, however, that this equivalence does not imply that the model-based approach and TD learning are identical across all scenarios. There has been a long debate in the reinforcement learning literature between model-based learning and TD learning, and we refer to the more relevant surveys and textbooks like Moerland et al. (2023) and Sutton and Barto (2018).
0:  The obtained trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$; Discount factor $\gamma$.
1:  $N(s,a)\leftarrow\sum_{i=1}^{T}\mathbb{I}\{s_{i}=s,a_{i}=a\}$, $N(s,a,s^{\prime})\leftarrow\sum_{i=1}^{T}\mathbb{I}\{s_{i}=s,a_{i}=a,s_{i+1}=s^{\prime}\}$, $\forall s,s^{\prime}\in{\mathcal{S}},a\in{\mathcal{A}}$
2:  $\widehat{r}(s)\leftarrow(\sum_{i=1}^{T}r_{i}\mathbb{I}\{s_{i}=s\})/(N(s)\vee 1)$, $\forall s\in{\mathcal{S}}$
3:  $\widehat{P}(s^{\prime}\mid s,a)=N(s,a,s^{\prime})/(N(s,a)\vee 1)$, $\forall s,s^{\prime}\in{\mathcal{S}},a\in{\mathcal{A}}$
4:  $\widehat{\bm{V}}^{\textnormal{MB}}\leftarrow({\bm{I}}-\gamma\widehat{\bm{P}})^{-1}\widehat{\bm{r}}$.
5:  return  $\widehat{\bm{V}}^{\textnormal{MB}}$.
Algorithm 1 Model-based Estimation
0:  The obtained trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$; Discount factor $\gamma$.
1:  Solve the following optimization problem Eq. (3),
$\min_{\widehat{\bm{V}}}\sum_{s\in{\mathcal{S}}}\left({\sum_{s_{i}=s}r(s_{i},a_{i})+\gamma\widehat{V}(s_{i+1})-\widehat{V}(s_{i})}\right)^{2}.$
2:  return  $\widehat{\bm{V}}^{\textnormal{TD}}$.
Algorithm 2 Model-free Estimation (Temporal Difference Learning)
3.1 Asymptotic Inference of Value Function
In this subsection, we investigate the asymptotic normality of the estimators $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$. For a fixed and static policy $\pi$, the MDP can induce a Markov chain where the states sequentially evolve. The states we observed on the trajectory $\tau$ are sampled from such a Markov chain, and the technical subtlety is that the samples are not necessarily independent and identically distributed (i.i.d.). Particularly, the current state $s_{t+1}$ depends on the last state $s_{t}$. However, if $s_{t}$ is sampled from the stationary distribution of the induced Markov chain, the future samples can be seen as i.i.d. generated from the stationary distribution. This observation allows us to hypothesize that rapid convergence to the stationary distribution may lead to asymptotic normality. To elucidate this concept further, we introduce the definition of uniform ergodicity, which indicates rapid mixing within the Markov chain. This principle is akin to those found in the works of Farias et al. (2022) and Hu and Wager (2022) and it is commonly assumed in the central limit theorem for Markov chains as detailed by Jones (2004).
Definition 3.3 (Uniform Ergodicity)
A Markov chain with the state transition matrix ${\bm{P}}$ is uniformly ergodic if it has stationary distribution ${\bm{\mu}}$ and there exists some constant $C\geq 0$ and $10\leq\lambda<$ 1 such that for all $k\in\mathbb{N}$ and all $s\in{\mathcal{S}}$,
$D_{\mathrm{TV}}({\bm{P}}^{k}(s,\cdot),{\bm{\mu}})\leq C\lambda^{k}\,,$
where $D_{\mathrm{TV}}(\cdot,\cdot)$ is the total variation distance.
Now we present the main theorem of this section, establishing the central limit theorem for $\widehat{\bm{V}}^{\textnormal{MB}}$, which naturally holds for $\widehat{\bm{V}}^{\textnormal{TD}}$. The detailed proof is delayed to Appendix 12.2.
Theorem 3.4 (Asymptotic Normality)
Assuming the Markov chain induced by policy $\pi$ is uniformly ergodic, with the input of a trajectory generated by a fixed policy $\pi$, it holds that as $T\to\infty$,
[…]
where
$\Sigma_{\mathrm{MB}}=({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-1}\operatorname{Diag}_{s\in{\mathcal{S}}}\left({\sigma^{2}_{\pi}(s)/\mu^{\pi}(s)+\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}/\mu^{\pi}(s)}\right)({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-\top}\,.$
We use $\sigma^{2}_{\pi}(s)\coloneqq\sum_{a\in{\mathcal{A}}}\mathbb{P}^{\pi}(a\mid s)\sigma^{2}(s,a)$ to denote the variance of the instant rewards of state $s$ under policy $\pi$. The vector ${\bm{\mu}}^{\pi}\coloneqq(\mu^{\pi}(s^{1}),\cdots,\mu^{\pi}(s^{K}))^{\top}$ is the stationary distribution under policy $\pi$, and $\Sigma_{p^{\pi}(s)}\coloneqq\operatorname{Diag}({\bm{p}}_{s})-{\bm{p}}_{s}{\bm{p}}_{s}^{\top}$, where ${\bm{p}}_{s}\in\mathbb{R}^{K}$ is the vector of the transition probability from state $s$ to the other states, $(\mathbb{P}^{\pi}(s^{1}\mid s),\cdots,\mathbb{P}^{\pi}(s^{K}\mid s))$.
Theorem 3.4 reveals that the model-based estimator is well-behaved and asymptotically normal. By Proposition 3.1, $\Sigma_{\mathrm{TD}}$ shares the same values and properties. In order to simplify and provide more insights, we discuss the inference of the value function under a specific state, say, the crucial state $s^{1}$. The following corollary presents the asymptotic normality of the estimator of $V(s^{1})$, which is the first dimension of ${\bm{V}}^{\textnormal{MB}}$ or ${\bm{V}}^{\textnormal{TD}}$.
Corollary 3.5
Following the same assumption as Theorem 3.4, it holds that as $T\to\infty$,
[…]
where
$\sigma^{2}_{\mathrm{MB}}(s^{1})=\sum_{s\in{\mathcal{S}}}\frac{\rho_{s^{1}}^{\pi}(s)^{2}}{\mu^{\pi}(s)}\left({\sigma_{\pi}^{2}(s)+\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}}\right)\,.$
Recall that $\rho^{\pi}_{s^{1}}(s)$ is the discounted state-visitation measure defined in Definition 2.1. The asymptotic variance of estimating $V(s^{1})$ is influenced not only by the state $s^{1}$ itself but also by all the other states. This is consistent with our understanding that our objective is the long-term value and the samples are drawn from an MDP. Although the variance depends on all states, we want to specially highlight that $\sigma^{2}_{\mathrm{MB}}(s^{1})$ can be decomposed into $K$ distinct terms, where each term corresponds almost exclusively to one state. Furthermore, each term is inversely proportion to the stationary probability $\mu^{\pi}(s)$. This implies that states which are less frequently visited pose greater challenges in sampling, potentially leading to increased variance. Additionally, each term has two parts, $\sigma_{\pi}^{2}(s)$ and $\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}$, which correspond to the (short-term) variance of the instant rewards and the (long-term) variance related to transitions to the other states.
3.2 Lower Bound for Asymptotic Variance of Estimating Value Function
In this subsection, we elaborate on the lower bound of the variance of any unbiased estimator. We reveal that the $\Sigma_{\mathrm{MB}}$ derived in Theorem 3.4 is optimal in the asymptotic view and cannot be further improved.
Theorem 3.6 (Asymptotic Efficiency)
@@ 3.2 Lower Bound for Asymptotic Variance of Estimating Value Function
Let the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ be sampled from MDP under policy $\pi$ with initial distribution ${\bm{\mu}}^{\pi}$. Then for any unbiased estimator of value functions with covariance matrix $\Sigma$, it holds that $\Sigma_{\mathrm{MB}}\preceq T\Sigma$, where $\Sigma_{\mathrm{MB}}$ is defined in Theorem 3.4.
To prove the theorem, let ${\bm{\theta}}=({\bm{P}}^{\pi},{\bm{r}}^{\pi})$ be the parameters controlling the system. Notice that ${\bm{\theta}}$ subjects to the constraint $\sum_{j}{\bm{P}}^{\pi}_{ij}=1$ for all $i\in[K]$. We then apply the Constrained Cramér–Rao bound (CCRB, see Lemma 18.5) for the lower bound. The full proof is delayed to Appendix 12.3.
In the recent literature, Khamaru et al. (2021) also establish a similar variance lower bound for TD learning, where they additionally assume the access to a generative model of the Markov chain. In particular, instead of sampling from a Markov chain, the generative model endows them with the flexibility to draw any random state transitions of the Markov chain. They can start from any state $s$ and generate the next state $s^{\prime}$ sampled from the distribution $\mathbb{P}^{\pi}(\cdot\mid s)$. With the generative model, they derive an asymptotic variance bound of TD as
$\Sigma_{\textnormal{GM}}=({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-1}\operatorname{Diag}_{s\in\mathcal{S}}\left({\sigma_{\pi}^{2}(s)+\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}}\right)({\bm{I}}-\gamma{\bm{P}}^{\pi})^{-\top}\,.$
Different from our $\Sigma_{\textnormal{MB}}$, such a lower bound shaves off the stationary distribution ${\bm{\mu}}^{\pi}$ at the denominators. This matches the intuition that we can only sample from the Markov chain instead of from the generative model.
@@ 4 A/B Testing on MDPs
In the previous section, we have established the inference results with the trajectory generated by a fixed policy $\pi$. To infer the ATE, however, it is necessary to simultaneously estimate ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ from a single trajectory. A commonly employed method for this purpose is A/B testing, where, at any given state $s$, a fair coin flip determines whether to apply $\pi^{\mathfrak{t}}$ or $\pi^{\mathfrak{c}}$. For simplicity, we introduce the experimental policy $\pi^{1/2}$, defined such that $\pi^{1/2}(a\mid s)\coloneqq\frac{1}{2}\cdot\pi^{\mathfrak{t}}(a\mid s)+\frac{1}{2}\cdot\pi^{\mathfrak{c}}(a\mid s)$ for any $(s,a)$ pair. Assume that the experimenter executes $\pi^{1/2}$ for $T$ epochs and obtains the trajectory $\tau=\left\{{(s_{i},a_{i},s_{i+1},r_{i})}\right\}_{i\in[T]}$. An example of a realized trajectory might include sequences such as:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s^{1},{\mathfrak{t}},s^{3},r_{1})}\quad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s^{3},{\mathfrak{c}},s^{3},r_{3})}\quad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s^{3},{\mathfrak{c}},s^{1},r_{3})}\quad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s^{1},{\mathfrak{c}},s^{2},r_{4})}\quad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s^{2},{\mathfrak{c}},s^{4},r_{5})}\quad{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s^{4},{\mathfrak{c}},s^{5},r_{6})}\cdots$
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\mathcal{D}}^{\mathfrak{t}}_{\mathrm{AB}}}=\{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(s_{i},a_{i},s_{i+1},r_{i})}:\pi_{i}=\pi^{\mathfrak{t}},i\in[T]\}\,,\qquad{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}{\mathcal{D}}^{\mathfrak{c}}_{\mathrm{AB}}}=\{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(s_{i},a_{i},s_{i+1},r_{i})}:\pi_{i}=\pi^{\mathfrak{c}},i\in[T]\}\,,$
where $\pi_{i}$ denotes policy used at time $i$. Leveraging the Markov property of the system, unbiased estimations of ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ can be derived from ${\mathcal{D}}^{\mathfrak{t}}$ and ${\mathcal{D}}^{\mathfrak{c}}$ respectively, using either Algorithm 1 or Algorithm 2. The corresponding algorithm is outlined in Algorithm 3.
@@
In the previous section, we have demonstrated that both Algorithm 1 and Algorithm 2 exhibit asymptotic normality. Given that these algorithms are directly employed in Algorithm 3, it should not be surprising to anticipate that A/B testing within the MDP framework may also achieve asymptotic normality. Formally, we have the following theorem.
@@
Assuming the Markov chain induced by the experiment policy $\pi^{1/2}$ is uniformly ergodic with stationary distribution ${\bm{\mu}}_{1/2}\coloneqq(\mu_{1/2}(s^{1}),\cdots,\mu_{1/2}(s^{K}))^{\top}$, it holds that as $T\to\infty$,
[…]
@@
$\Sigma_{\mathrm{AB}}$ $=2({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{t}})^{-1}\operatorname{Diag}_{s\in{\mathcal{S}}}\left({\frac{\sigma_{\mathfrak{t}}^{2}(s)}{\mu_{1/2}(s)}+\frac{\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s)}^{2}}{\mu_{1/2}(s)}}\right)({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{t}})^{-\top}$
$\qquad+2({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{c}})^{-1}\operatorname{Diag}_{s\in{\mathcal{S}}}\left({\frac{\sigma_{\mathfrak{c}}^{2}(s)}{\mu_{1/2}(s)}+\frac{\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}{\mu_{1/2}(s)}}\right)({\bm{I}}-\gamma{\bm{P}}^{\mathfrak{c}})^{-\top}\,.$
Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
[…]
where
$\sigma_{\mathrm{AB}}^{2}(s^{1})=2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\Big{(}\sigma_{\mathfrak{t}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}}(s)}^{2}\Big{)}+2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\left({\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}}(s)}^{2}}\right)\,.$
Although similar to Theorem 3.4 and Corollary 3.5, the terms in $\Sigma_{\mathrm{AB}}$ and $\sigma_{\mathrm{AB}}^{2}(s^{1})$ in A/B testing are split into two distinct components: one associated exclusively with $\pi^{\mathfrak{t}}$ and the other with $\pi^{\mathfrak{c}}$. Each term in $\Sigma_{\mathrm{AB}}$ and $\sigma_{\mathrm{AB}}^{2}(s^{1})$ includes an additional constant factor of 2 compared to $\Sigma_{\mathrm{MB}}$ and $\sigma_{\mathrm{MB}}^{2}(s^{1})$ outlined in Theorem 3.4 and Corollary 3.5. This reflects the fact that A/B testing employs a mixed policy, with $\pi^{\mathfrak{t}}$ and $\pi^{\mathfrak{c}}$ each applied half of the time in expectation. Furthermore, the denominator features $\mu_{1/2}(s)$ rather than $\mu^{\pi^{\mathfrak{t}}}(s)$ or $\mu^{\pi^{\mathfrak{c}}}(s)$, indicating that the stationary distribution is linked to the mixed policy $\pi^{1/2}$.
@@
4.1 Lower Bound for Experimentation with General Treatment

### v3

@@ 2.1 Preliminaries: Inference under a Fixed Policy
The two inference algorithms are refereed as Algorithm 2 and 3. Despite of the difference in nature, $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ are actually providing the same point estimation of the value function $\bm{V}$} under very mild assumptions.
@@
Let $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ be the respective output of Algorithm 2 and 3 given the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ sampled under policy $\pi$. Assume all states are visited at least once in $\tau$, then it holds $\widehat{\bm{V}}^{\textnormal{MB}}=\widehat{\bm{V}}^{\textnormal{TD}}$.
@@
With such an equivalence, establishing a property for one estimator inherently confirms it for the other. It is important to emphasize that this equivalence does not imply that the model-based approach and TD learning are identical across all scenarios. We are only talking about finite-space and finite-action MDPs without function approximations in this work. There has been a long debate in the RL literature between model-based learning and TD learning, and we refer to the more relevant surveys and textbooks like Moerland et al. (2023) and Sutton and Barto (2018).
Under mild technical conditions, $\widehat{\bm{V}}^{\textnormal{MB}}$ and $\widehat{\bm{V}}^{\textnormal{TD}}$ can be shown to enjoy asymptotic normality.
Proposition 2.2
Assume the Markov chain induced by a fixed policy $\pi$ is uniformly ergodic. Given a trajectory under any initial distribution, it holds as $T\to\infty$, $\sqrt{T}(\widehat{\bm{V}}^{\textnormal{MB}}-{\bm{V}}^{\pi})\xrightarrow{d}{\mathcal{N}}({\bm{0}},\Sigma_{\mathrm{MB}})$.
Here uniform ergodicity is a common technical assumption also used in Farias et al. (2022), Hu and Wager (2022), Jones (2004) and will be specified in Appendix 12.1 along with the detailed form of $\Sigma_{\mathrm{MB}}$. While we expected this result to have been established previously in the RL literature, we were unable to locate an exact reference. For completeness, we provide a full proof in Appendix 12.1. In addition, we highlight another important result, the asymptotic efficiency of these two classical estimators, which may be of independent interest to the RL literature. Specifically, $\Sigma_{\mathrm{MB}}$ is essentially the smallest variance that can be obtained by any unbiased estimator. The proof is given in Appendix 13.1.
Proposition 2.3
@@
Let the trajectory $\tau=\{(s_{i},a_{i},s_{i+1},r_{i})\}_{i\in[T]}$ be sampled from MDP under policy $\pi$ with the stationary distribution ${\bm{\mu}}^{\pi}$. Then for any unbiased estimator of value functions with covariance matrix $\Sigma$, it holds that $\Sigma_{\mathrm{MB}}\preceq T\Sigma$.
@@ 3 A/B Testing on MDPs
In the previous section, we have established the inference results with the trajectory generated by a fixed policy $\pi$. To infer the ATE, however, it is necessary to simultaneously estimate ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ from a single trajectory. A commonly employed method for this purpose is A/B testing, where, at any given state $s$, a fair coin flip determines whether to apply $\pi^{\mathfrak{t}}$ or $\pi^{\mathfrak{c}}$. For simplicity, we introduce the experimental policy $\pi^{1/2}$, defined such that $\pi^{1/2}(a\mid s)\coloneqq{1}/{2}\cdot\pi^{\mathfrak{t}}(a\mid s)+{1}/{2}\cdot\pi^{\mathfrak{c}}(a\mid s)$ for any $(s,a)$ pair. Assume that the experimenter executes $\pi^{1/2}$ for $T$ epochs and obtains the trajectory $\tau=\left\{{(s_{i},a_{i},s_{i+1},r_{i})}\right\}_{i\in[T]}$. An example of a realized trajectory might include sequences such as:
${\color[rgb]{0,0,1}(s^{1},{\mathfrak{t}},s^{3},r_{1})}\quad{\color[rgb]{1,.5,0}(s^{3},{\mathfrak{c}},s^{3},r_{2})}\quad{\color[rgb]{0,0,1}(s^{3},{\mathfrak{c}},s^{1},r_{3})}\quad{\color[rgb]{1,.5,0}(s^{1},{\mathfrak{c}},s^{2},r_{4})}\quad{\color[rgb]{1,.5,0}(s^{2},{\mathfrak{c}},s^{4},r_{5})}\quad{\color[rgb]{0,0,1}(s^{4},{\mathfrak{c}},s^{5},r_{6})}\cdots$
@@
${\color[rgb]{0,0,1}{\mathcal{D}}^{\mathfrak{t}}_{\mathrm{AB}}}=\{{\color[rgb]{0,0,1}(s_{i},a_{i},s_{i+1},r_{i})}:\pi_{i}=\pi^{\mathfrak{t}},i\in[T]\}\,,\qquad{\color[rgb]{1,.5,0}{\mathcal{D}}^{\mathfrak{c}}_{\mathrm{AB}}}=\{{\color[rgb]{1,.5,0}(s_{i},a_{i},s_{i+1},r_{i})}:\pi_{i}=\pi^{\mathfrak{c}},i\in[T]\}\,$,
where $\pi_{i}$ denotes policy used at time $i$. Leveraging the Markov property of the system, unbiased estimations of ${\bm{V}}^{\mathfrak{t}}$ and ${\bm{V}}^{\mathfrak{c}}$ can be derived from ${\mathcal{D}}^{\mathfrak{t}}$ and ${\mathcal{D}}^{\mathfrak{c}}$ respectively, using either Algorithm 2 or Algorithm 3. The corresponding algorithm is outlined in Algorithm 1. Although the idea is straightforwards, the estimator can perform in our desired way. More specifically, we can establish the asymptotic normality.
@@
Assuming the Markov chain induced by the experiment policy $\pi^{1/2}$ is uniformly ergodic, it holds that $\sqrt{T}\left({\widehat{\bm{\Delta}}^{\mathrm{AB}}-{\bm{\Delta}}}\right)\xrightarrow{d}{\mathcal{N}}({\bm{0}},\Sigma_{\mathrm{AB}})$ as $T\to\infty$. The detailed form of $\Sigma_{\mathrm{AB}}$ is provided in the appendix (Theorem 12.4).
Specifically, for the crucial state $s^{1}$, it holds that as $T\to\infty$,
$\sqrt{T}\left({\widehat{\Delta}^{\mathrm{AB}}(s^{1})-\Delta(s^{1})}\right)\xrightarrow{d}{\mathcal{N}}(0,\sigma_{\mathrm{AB}}^{2}(s^{1}))\,$,
@@
$\sigma_{\mathrm{AB}}^{2}(s^{1})=2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{t}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\Big{(}\sigma_{\mathfrak{t}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{t}}}\right\|_{\Sigma_{p^{\mathfrak{t}}(s)}}^{2}\Big{)}+2\sum_{s\in{\mathcal{S}}}\frac{\rho^{\mathfrak{c}}_{s^{1}}(s)^{2}}{\mu_{1/2}(s)}\left({\sigma_{\mathfrak{c}}^{2}(s)+\left\|{\gamma{\bm{V}}^{\mathfrak{c}}}\right\|_{\Sigma_{p^{\mathfrak{c}}(s)}}^{2}}\right)\,$,
where $\rho^{\pi}_{s^{1}}(s)$ is defined as $\sum_{i=0}^{\infty}\gamma^{i}\mathbb{P}^{\pi}(s_{i}=s\mid s_{0}=s^{1})$. We use $\sigma^{2}_{\pi}(s)\coloneqq\sum_{a\in{\mathcal{A}}}\mathbb{P}^{\pi}(a\mid s)\sigma^{2}(s,a)$ to denote the variance of the instant rewards of state $s$ under policy $\pi$. The term $\mu_{1/2}(s)$ is the stationary probability of state $s$ under the experiment policy $\pi^{1/2}$, and $\Sigma_{p^{\pi}(s)}\coloneqq\operatorname{Diag}({\bm{p}}_{s})-{\bm{p}}_{s}{\bm{p}}_{s}^{\top}$, where ${\bm{p}}_{s}\in\mathbb{R}^{K}$ is the vector of the transition probability from state $s$ to the other states, $(\mathbb{P}^{\pi}(s^{1}\mid s),\cdots,\mathbb{P}^{\pi}(s^{K}\mid s))^{\top}$.
The term $\rho^{\pi}_{s^{1}}(s)$ is also called the discounted state-visitation measure defined and widely used in the RL literature (Agarwal et al. 2022). The asymptotic variance $\sigma_{\mathrm{AB}}^{2}(s^{1})$ is influenced not only by the state $s^{1}$ itself but also by all the other states. This is consistent with our understanding that our objective is the long-term value and the samples are drawn from an MDP. Although the variance depends on all states, we want to specially highlight that $\sigma^{2}_{\mathrm{AB}}(s^{1})$ can be divided into two groups: one associated exclusively with $\pi^{\mathfrak{t}}$ and the other with $\pi^{\mathfrak{c}}$. Each group has $|{\mathcal{S}}|$ distinct terms, where each term corresponds almost exclusively to one state. Furthermore, each term is inversely proportion to the stationary probability $\mu_{1/2}(s)$. This implies that states which are less frequently visited pose greater challenges in sampling, potentially leading to increased variance. Additionally, each term has two parts, $\sigma_{\pi}^{2}(s)$ and $\left\|{\gamma{\bm{V}}^{\pi}}\right\|_{\Sigma_{p^{\pi}(s)}}^{2}$, which correspond to the (short-term) variance of the instant rewards and the (long-term) variance related to transitions to the other states.
@@
3.1 Variance Lower Bound for General A/B Testing on MDPs